# SAGA: Source Attribution of Generative AI Videos

### v1

@@ Abstract
The proliferation of generative AI has led to hyper-realistic synthetic videos, escalating misuse risks and outstripping binary real/fake detectors. We introduce SAGA (Source Attribution of Generative AI videos), the first comprehensive framework to address the urgent need for AI-generated video source attribution at a large scale. Unlike traditional detection, SAGA identifies the specific generative model used. It uniquely provides multi-granular attribution across five levels: authenticity, generation task (e.g., T2V/I2V), model version, development team, and the precise generator, offering far richer forensic insights. Our novel video transformer architecture, leveraging features from a robust vision foundation model, effectively captures spatio-temporal artifacts. Critically, we introduce a data-efficient pretrain-and-attribute strategy, enabling SAGA to achieve state-of-the-art attribution using only 0.5% of source-labeled data per class, matching fully supervised performance. Furthermore, we propose Temporal Attention Signatures (T-Sigs), a novel interpretability method that visualizes learned temporal differences, offering the first explanation for why different video generators are distinguishable. Extensive experiments on public datasets, including cross-domain scenarios, demonstrate that SAGA sets a new benchmark for synthetic video provenance, providing crucial, interpretable insights for forensic and regulatory applications.
@@ 1 Introduction
The rapid advancement of AI-driven video synthesis, spanning text-to-video (T2V) [wang2023modelscope, opensora, morphstudio] and image-to-video (I2V) [blattmann2023stable, xing2024dynamicrafter, chen2023seine] systems, has democratized content creation but also heightened concerns over misuse and misinformation [sora2024, lin2025omnihuman], exemplified by incidents like AI-generated wildfire videos causing public alarm [reuters2025wildfires]. Current defenses largely focus on binary real/fake detection [chen2024demamba, kundu2025towards, xu2023tall, cheng2024can]. However, as generative models multiply and evolve at an unprecedented pace [li2024fakebench, lee2024tug], merely detecting a video as synthetic is insufficient. The critical need has shifted from whether it’s fake to what is its source? Identifying the specific generative model or family, called source attribution [wang2023did, yang2022deepfake, vahdati2024beyond], is paramount for effective digital forensics [amerini2024deepfake], intellectual property enforcement [ballardini2019ai, smits2022generative], and developing robust adversarial countermeasures [pinhasov2024xai, zhou2022adversarial].
@@
Attributing synthetic videos to their source is a far more complex challenge than traditional DeepFake detection or even image source attribution [wang2023did, yang2022deepfake, fang2023open]. While image-based methods offer a starting point, they fundamentally fail to address video-specific complexities. We identify three key barriers: (1) Temporal Dynamics: Videos possess unique temporal fingerprints and inconsistencies resulting from the generation process, entirely missed by static image analysis. (2) Increased Model Diversity: The video generation pipeline involves more diverse architectures and stages (e.g., frame synthesis, motion models), creating a vastly larger and more complex attribution space. (3) Video Compression: Unlike image compression, video codecs introduce complex spatio-temporal artifacts that can obscure or destroy subtle generator-specific traces. These challenges necessitate a novel approach designed specifically for the video domain.
@@
To address this significant gap, we introduce SAGA (Source Attribution of Generative AI videos), the first large-scale, comprehensive framework dedicated to multi-granular source attribution of AI-generated videos. Moving decisively beyond binary detection, SAGA pinpoints the origin of a synthetic video across five crucial levels of granularity using only 0.5% of the data: (1) BIN-L (real/synthetic); (2) TASK-L (real vs. T2V vs. I2V); (3) SD-L (differentiating between Stable Diffusion versions e.g., [rombach2022high, podell2023sdxl]); (4) TEAM-L (attributing to development teams, aiding misuse tracking); and (5) GEN-L (precise model ID). This multi-granular approach is crucial in practice: for example, when two generators are highly similar, an in-the-wild video may yield low-confidence predictions at the GEN-L, but higher confidence at coarser levels such as SD version or team, still providing valuable forensic insight. Furthermore, unlike prior works, SAGA provides Temporal Attention Signatures or T-Sig as shown in Fig. 1(a), which offer crucial interpretability into why generative models are distinguishable.
@@
By averaging frame-to-frame attention scores across multiple videos from a common source, we derive unique visual ‘fingerprints’ (T-Sig) for each generator. To the best of our knowledge, this is the first work to visually explain video attribution performance: T-Sig highlight the subtle but stable temporal artifacts, such as characteristic motion dynamics or frame-to-frame inconsistencies, that SAGA learns to leverage for fine-grained source identification.
@@
To achieve this, SAGA employs a novel multi-headed attention video transformer to effectively capture temporal inconsistencies. To enhance in-the-wild robustness, we initialize our model with rich visual features from a foundational vision encoder [alabdulmohsin2024getting], mitigating domain gap issues [chen2025finger, lv2024domainforensics, chen2021featuretransfer]. Addressing the common scenario of abundant binary labels but scarce multi-class source labels, we propose a pretrain-and-adapt strategy. We first build a strong visual representation by pretraining a binary (real vs. fake) classifier. Subsequently, this base model is efficiently adapted to the multi-class source attribution challenge, utilizing a contrastive objective with hard-negative mining (HNM). Remarkably, this adaptation, allows SAGA to match the performance of a model trained with 100% of the source-labeled data, even when using only 0.5% of the labeled examples for the adaptation phase (Fig. 1(c)). This highlights exceptional data efficiency for the complex task of synthetic video source attribution.
Figure 2: Overall framework of SAGA with a two-stage training approach. In Stage-1, each video $x_{k}$ with real/fake labels is processed through a frozen foundational vision encoder to extract image-level features $z_{m}$, which are stacked in temporal order to form the video representation $\zeta_{k}$. Positional encoding is added, and the sequence is passed through our video transformer architecture $\theta$ (Sec. 3.1) to obtain $\phi_{k}$. The classifier $\beta_{1}$ maps $\phi_{k}$ to real or fake classes using a cross-entropy loss ($\mathcal{L}_{CE}$). In Stage-2, the pretrained video transformer is adapted for attribution into $n_{c}$ classes ($n_{c}$ defined by the attribution task; see Sec. 3) using only 0.5% of source labeled data. Stage-2 incorporates an additional hard negative mining objective ($\mathcal{L}_{\text{HNM}}$, Sec. 3.2) along with $\mathcal{L}_{CE}$ for the attribution task.
@@
  • We conduct extensive evaluations across 19 distinct video generators from two public datasets (DeMamba [chen2024demamba] and DVF [song2024dvfdataset]). Our results validate SAGA’s effectiveness and robustness in multi-granular source attribution under both in-domain and cross-domain scenarios, setting a new benchmark for this emerging field.
@@ 2 Related Work
Synthetic Content Detection: Early efforts in synthetic content detection primarily targeted images, a focus driven by the evolution of generative models, from early GAN models [karras2020analyzing, karras2019style, zhu2017unpaired] to the more recent and powerful diffusion models [rombach2022high, ramesh2022hierarchical, saharia2022photorealistic]. Corvi et al. [corvi2023detection, corvi2023intriguing] investigated the detection of diffusion-generated images from GAN detectors through spatial and frequency domain analyses, and revealed the distinctive forensic traces left by generation models. Several studies explored diverse features to distinguish realistic diffusion-generated images, including reconstruction errors[wang2023dire], CLIP-based representations [ojha2023towards, tan2025c2p], and up-sampling artifacts [tan2024rethinking]. However, as demonstrated in [chen2024demamba, vahdati2024beyond], image-centric approaches fall short when applied to videos, where the capture of different spatial traces or temporal artifacts is essential for effective detection. Synthetic video detection remains comparatively underexplored but has gained more focus in the recent two years. DeMamba [chen2024demamba] addresses this by employing a structured state space model that continuously scans spatial and temporal zones to capture subtle generative artifacts, enabling robust real/fake classification across diverse video generators and outperforming image-based detectors on video-specific inconsistencies. A large AI-generated video dataset with 19 T2V/I2V generators is also proposed. UNITE [kundu2025towards] tackles face-manipulation and synthetic video detection using a foundation model with a transformer and attention-diversity loss. While these methods advance video authenticity detection, they are primarily confined to binary classification and do not address the more challenging task of source attribution.
@@
Source Model Attribution: Research in source attribution has been largely focused on synthetic images. Girish et al. [girish2021towards] proposed an open-world discovery and attribution pipeline that iteratively combines out-of-distribution detection, clustering, and supervised refinement, enabling the discovery and attribution of images from both known and unknown GANs in a scalable manner. POSE [yang2023progressive] further advanced this by simulating open-set samples using lightweight augmentation models to better model the imperceptible traces left by unknown generative models. Wang et al. [wang2023did] tackled the origin attribution problem from a model-agnostic and alteration-free perspective, proposing a reverse-engineering approach that leverages reconstruction loss: if an image can be more accurately inverted by a given model, it is likely to have been generated by that model. Collectively, these works highlight the shift from simple real/fake detection to fine-grained, open-set, and model-agnostic attribution in images. However, most of these methods are tailored for static images and cannot address the unique spatiotemporal challenges of source attribution in synthetic videos, where temporal consistency and motion artifacts play a critical role.
@@
To the best of our knowledge, the only prior work to attempt source attribution in videos is by Vahdati et al. [vahdati2024beyond], whose study is limited to only 4 generators with closed-source videos and focuses on only generator-level attribution. In contrast, SAGA provides a far more comprehensive benchmark (as shown in Table 1) on 19 video generators, spanning multiple levels of attribution granularity along with interpretable T-Sig analyses for actionable provenance analysis.
@@ 3.1 Video Transformer
AI-generated videos inherently exhibit domain gap [chen2025finger, chen2021featuretransfer, lv2024domainforensics], which is critical to address since the aim of SAGA is to be used in-the-wild. To enhance robustness, we extract potentially domain-agnostic features by leveraging a powerful visual encoder pretrained on web-scale image-text data. Given a video instance $x_{k}\in\mathcal{X}$, we process each frame $g_{m}$ (resized to a fixed resolution) using the frozen pretrained encoder. This produces a tokenized embedding $z_{m}\in\mathbb{R}^{l_{t}\times d_{t}}$ for each frame, where $m\in{1,2,\ldots,L}$ with $L$ denoting the number of frames per video. The dimension of these embeddings is determined by the chosen encoder, where $l_{t}$ is the number of tokens per frame and $d_{t}$ is the token feature dimension. The embeddings for all frames in $x_{k}$ are concatenated in temporal order, resulting in a video-level representation $\zeta_{k}\in\mathbb{R}^{L\times l_{t}\times d_{t}}$, which serves as input to our trainable video transformer. The resulting set of encoded videos is thus represented as $\mathcal{Z}={\zeta_{k}\mid x_{k}\in\mathcal{X}}$. The SAGA’s video-transformer model employs a multi-head self-attention (MHSA) transformer architecture [vaswani2017attention] ($\theta$) tailored for video attribution to obtain $\phi_{k}=\theta(\zeta_{k})$. It processes sequences of frame embeddings, effectively capturing temporal dependencies for robust video-level predictions.
@@
  • A two-layer feed-forward network (MLP) with GELU activation [hendrycks2016gaussian] for non-linear transformations.
@@ 3.2 Contrastive Objective
Semi-hard negatives, which are most commonly used in the literature [schroff2015facenet, harwood2017smart, xuan2020improved], are those that are further from the anchor than the positive, but within the margin as,
@@
Table 5: SOTA and SAGA comparison on DeMamba [chen2024demamba] for in-domain and DVF [song2024dvfdataset] for cross-domain evaluation. Best and second-best performances are highlighted.
@@
DeMamba [chen2024demamba] TALL [xu2023tall] 87.91% 88.52% 88.42%
F3Net [qian2020thinking] 88.73% 81.88% 86.04%
NPR [tan2024rethinking] 82.45% 84.08% 83.45%
STIL [gu2021spatiotemporal] 87.12% 82.22% 85.35%
MINTIME-CLIP-B [chen2024demamba] 91.55% 87.62% 89.98%
FTCN-CLIP-B [chen2024demamba] 92.21% 86.18% 89.67%
CLIP-B-PT [chen2024demamba] 44.83% 81.74% 41.82%
DeMamba-CLIP-PT [chen2024demamba] 79.97% 78.86% 79.98%
XCLIP-B-PT [chen2024demamba] 61.29% 81.93% 65.83%
DeMamba-XCLIP-PT [chen2024demamba] 76.38% 83.59% 79.31%
XCLIP-B-FT [chen2024demamba] 86.77% 84.41% 86.07%
@@
DVF [song2024dvfdataset] CNNDet [wang2020cnn] - - 78.20%
DIRE [wang2023dire] - - 62.10%
Raising [cozzolino2024raising] - - 67.00%
UNI-FD [ojha2023towards] - - 74.10%
F3Net [qian2020thinking] - - 81.30%
ViViT [arnab2021vivit] - - 79.10%
TALL [xu2023tall] - - 69.50%
TS2-Net [liu2022ts2] - - 72.10%
DE-FAKE [sha2023fake] - - 72.10%
HifiNet [guo2023hierarchical] - - 84.30%
DVF [song2024dvfdataset] - - 92.00%
@@ 4 Experiments
Datasets: The training is performed on the DeMamba [chen2024demamba] dataset, with 19 different AI video generators and 1M real videos. More details can be found in the supplementary material, including how we define different attribution levels. Additionally, we use the DVF dataset [song2024dvfdataset] covering 8 video generators for cross-data evaluations. Implementation details are provided in the supplementary.
Real/Fake Detection Results: We first comprehensively evaluate the performance of SAGA in binary classification in Table 4, conducting both in-domain and cross-generator evaluations. For cross-generator analysis, we train on the train split of DeMamba and evaluate on the val split, as well as train on generators from a specific team or SD version backbone or generation task and test on the remaining generators. The results demonstrate that, for the BIN-L task, SAGA achieves robust authenticity verification across diverse and previously unseen data sources, exhibiting strong generalization and minimal sensitivity to domain shifts. To rigorously benchmark SAGA against existing state-of-the-art (SOTA) binary detectors, we conducted evaluations on both the DeMamba [chen2024demamba] and DVF [song2024dvfdataset] datasets, as presented in Table 5. Notably, all competing SOTA methods on the DVF dataset [song2024dvfdataset] are trained and tested within the DVF dataset, following standard in-domain protocols. In contrast, SAGA is trained exclusively on DeMamba and evaluated directly on DVF, constituting a challenging cross-dataset generalization scenario. Despite this, SAGA significantly outperforms the SOTA baselines, highlighting its superior robustness and generalizability for the BIN-L task across diverse data distributions.
@@
t-SNE Analysis: To further interpret the representations learned by SAGA, we conduct a t-SNE [van2008visualizing] and T-Sig analyses of the feature embeddings produced by the model under different attribution settings. Specifically, we visualize the embedding outputs of the last ($6^{th}$) encoder of $\theta$ for videos from the validation set when the model is trained for all attribution levels. First, Fig. 4 (a) and (b) with TASK-L attribution results show that embeddings for real, T2V, and I2V samples form clearly separable clusters, demonstrating effective discrimination among these broad categories. However, when the same embeddings are colored by generator, substantial overlap is observed among most generators, with only a few, such as MorphStudio [morphstudio] and SVD [blattmann2023stable], forming distinct clusters. This suggests that while SAGA is highly effective at coarse-grained attribution, it does not inherently separate individual generators at this level. Fig. 4 (c) shows the t-SNE plot for the SAGA model trained for BIN-L classification. Here, all generators except Pika [pika2022] collapse into a single “fake” cluster, indicating that the model learns to aggregate all synthetic sources together for the binary task, with minimal separation among generators.
Fig. 4 (d) visualizes embeddings from a SAGA model trained on 10 specific generators from the DeMamba train set. The seen generators form distinct clusters, and notably, several unseen generators such as Hotshot [hotshot2023], Show_1 [zhang2024show], and MorphStudio [morphstudio] also appear as separable clusters. This indicates that the model, even trained on a subset of generators, can recognize distributional differences and cluster unseen sources, highlighting its potential for generalization in open-set scenarios. For the SD-L and TEAM-L models, t-SNE projections (Fig. 4 (e) - (h)) reveal that the learned representations not only cluster according to the supervised SD version or team labels, but also often separate individual generators within each group. This indicates that SAGA captures fine-grained differences between generators, even when supervision is provided only at a coarser level. This level of separation indicates that the model is sensitive to subtle distributional differences introduced by specific generator architectures or research teams, enabling it to infer whether an unknown generator shares an SD backbone or team affiliation, or represents a completely novel source.

### v2

@@ Abstract
The proliferation of generative AI has led to hyper-realistic synthetic videos, escalating misuse risks and outstripping binary real/fake detectors. We introduce SAGA (Source Attribution of Generative AI videos), the first comprehensive framework to address the urgent need for AI-generated video source attribution at a large scale. Unlike traditional detection, SAGA identifies the specific generative model used. It uniquely provides multi-granular attribution across five levels: authenticity, generation task (e.g., T2V/I2V), model version, development team, and the precise generator, offering far richer forensic insights. Our novel video transformer architecture, leveraging features from a robust vision foundation model, effectively captures spatio-temporal artifacts. Critically, we introduce a data-efficient pretrain-and-attribute strategy, enabling SAGA to achieve state-of-the-art attribution using only 0.5% of source-labeled data per class, matching fully supervised performance. Furthermore, we propose Temporal Attention Signatures (T-Sigs), a novel interpretability method that visualizes learned temporal differences, offering the first explanation for why different video generators are distinguishable. Extensive experiments on public datasets, including cross-domain scenarios, demonstrate that SAGA sets a new benchmark for synthetic video provenance, providing crucial, interpretable insights for forensic and regulatory applications. The project page is https://rohit-kundu.github.io/SAGA.
@@ 1 Introduction
The rapid advancement of AI-driven video synthesis, spanning text-to-video (T2V) [49, 59, 43] and image-to-video (I2V) [5, 53, 9] systems, has democratized content creation but also heightened concerns over misuse and misinformation [31, 26], exemplified by incidents like AI-generated wildfire videos causing public alarm [36]. Current defenses largely focus on binary real/fake detection [7, 22, 54, 10]. However, as generative models multiply and evolve at an unprecedented pace [25, 24], merely detecting a video as synthetic is insufficient. The critical need has shifted from whether it’s fake to what is its source? Identifying the specific generative model or family, called source attribution [52, 56, 46], is paramount for effective digital forensics [2], intellectual property enforcement [4, 41], and developing robust adversarial countermeasures [32, 60].
@@
Attributing synthetic videos to their source is a far more complex challenge than traditional DeepFake detection or even image source attribution [52, 56, 14]. While image-based methods offer a starting point, they fundamentally fail to address video-specific complexities. We identify three key barriers: (1) Temporal Dynamics: Videos possess unique temporal fingerprints and inconsistencies resulting from the generation process, entirely missed by static image analysis. (2) Increased Model Diversity: The video generation pipeline involves more diverse architectures and stages (e.g., frame synthesis, motion models), creating a vastly larger and more complex attribution space. (3) Video Compression: Unlike image compression, video codecs introduce complex spatio-temporal artifacts that can obscure or destroy subtle generator-specific traces. These challenges necessitate a novel approach designed specifically for the video domain.
@@
To address this significant gap, we introduce SAGA (Source Attribution of Generative AI videos), the first large-scale, comprehensive framework dedicated to multi-granular source attribution of AI-generated videos. Moving decisively beyond binary detection, SAGA pinpoints the origin of a synthetic video across five crucial levels (denoted “-L”) of granularity using only 0.5% of the data: (1) BIN-L (real/synthetic); (2) TASK-L (real vs. T2V vs. I2V); (3) SD-L (differentiating between Stable Diffusion versions e.g., [37, 33]); (4) TEAM-L (attributing to development teams, aiding misuse tracking); and (5) GEN-L (precise model ID). This multi-granular approach is crucial in practice: for example, when two generators are highly similar, an in-the-wild video may yield low-confidence predictions at the GEN-L, but higher confidence at coarser levels such as SD version or team, still providing valuable forensic insight. Furthermore, unlike prior works, SAGA provides Temporal Attention Signatures or T-Sig as shown in Fig. 1(a), which offer crucial interpretability into why generative models are distinguishable.
@@
By averaging frame-to-frame attention scores across multiple videos from a common source, we derive unique visual ‘fingerprints’ (T-Sig) for each generator. To the best of our knowledge, this is the first work to visually explain video attribution performance: T-Sigs highlight the subtle but stable temporal artifacts, such as characteristic motion dynamics or frame-to-frame inconsistencies, that SAGA learns to leverage for fine-grained source identification.
@@
To achieve this, SAGA employs a novel multi-headed attention video transformer to effectively capture temporal inconsistencies. To enhance in-the-wild robustness, we initialize our model with rich visual features from a foundational vision encoder [1], mitigating domain gap issues [8, 28, 6]. Addressing the common scenario of abundant binary labels but scarce multi-class source labels, we propose a pretrain-and-adapt strategy. We first build a strong visual representation by pretraining a binary (real vs. fake) classifier. Subsequently, this base model is efficiently adapted to the multi-class source attribution challenge, utilizing a contrastive objective with hard-negative mining (HNM). Remarkably, this adaptation, allows SAGA to match the performance of a model trained with 100% of the source-labeled data, even when using only 0.5% of the labeled examples for the adaptation phase (Fig. 1(c)). This highlights exceptional data efficiency for the complex task of synthetic video source attribution.
Figure 2: Overall framework of SAGA with a two-stage training approach. In Stage-1, each video $x_{k}$ with real/fake labels is processed through a frozen foundational vision encoder to extract image-level features $z_{m}$, which are stacked in temporal order to form the video representation $\zeta_{k}$. Positional encoding is added, and the sequence is passed through our video transformer architecture $\theta$ (Sec. 3.1) to obtain $\phi_{k}$. The classifier $\beta_{1}$ maps $\phi_{k}$ to real or fake classes using a cross-entropy loss ($\mathcal{L}_{CE}$). In Stage-2, the pretrained video transformer is adapted for attribution into $n_{c}$ classes ($n_{c}$ defined by the attribution task; see supplementary Sec. S-1) using only 0.5% of source labeled data. Stage-2 incorporates an additional hard negative mining objective ($\mathcal{L}_{\text{HNM}}$, Sec. 3.2) along with $\mathcal{L}_{CE}$ for the attribution task.
@@
  • We conduct extensive evaluations across 19 distinct video generators from two public datasets (DeMamba [7] and DVF [42]). Our results validate SAGA’s effectiveness and robustness in multi-granular source attribution under both in-domain and cross-domain scenarios, setting a new benchmark for this emerging field.
@@ 2 Related Work
Synthetic Content Detection: Early efforts in synthetic content detection primarily targeted images, a focus driven by the evolution of generative models, from early GAN models [21, 20, 61] to the more recent and powerful diffusion models [37, 35, 38]. The authors in [12, 11] investigated the detection of diffusion-generated images from GAN detectors through spatial and frequency domain analyses, and revealed the distinctive forensic traces left by generation models. Several studies explored diverse features to distinguish realistic diffusion-generated images, including reconstruction errors[51], CLIP-based representations [30, 44], and up-sampling artifacts [45]. However, as demonstrated in [7, 46], image-centric approaches fall short when applied to videos, where the capture of different spatial traces or temporal artifacts is essential for effective detection. Synthetic video detection remains comparatively underexplored but has gained more focus in the recent two years. DeMamba [7] addresses this by employing a structured state space model that continuously scans spatial and temporal zones to capture subtle generative artifacts, enabling robust real/fake classification across diverse video generators and outperforming image-based detectors on video-specific inconsistencies. A large AI-generated video dataset with 19 T2V/I2V generators is also proposed. UNITE [22] tackles face-manipulation and synthetic video detection using a foundation model with a transformer and attention-diversity loss. While these methods advance video authenticity detection, they are primarily confined to binary classification and do not address the more challenging task of source attribution.
@@
Source Model Attribution: Research in source attribution has been largely focused on synthetic images. [15] proposed an open-world discovery and attribution pipeline that iteratively combines out-of-distribution detection, clustering, and supervised refinement, enabling the discovery and attribution of images from both known and unknown GANs in a scalable manner. POSE [57] further advanced this by simulating open-set samples using lightweight augmentation models to better model the imperceptible traces left by unknown generative models. Wang et al. [52] tackled the origin attribution problem from a model-agnostic and alteration-free perspective, proposing a reverse-engineering approach that leverages reconstruction loss: if an image can be more accurately inverted by a given model, it is likely to have been generated by that model. Collectively, these works highlight the shift from simple real/fake detection to fine-grained, open-set, and model-agnostic attribution in images. However, most of these methods are tailored for static images and cannot address the unique spatiotemporal challenges of source attribution in synthetic videos, where temporal consistency and motion artifacts play a critical role.
@@
To the best of our knowledge, the only prior work to attempt source attribution in videos is by Vahdati et al. [46], whose study is limited to only 4 generators with closed-source videos and focuses on only generator-level attribution. In contrast, SAGA provides a far more comprehensive benchmark (as shown in Table 1) on 19 video generators, spanning multiple levels of attribution granularity along with interpretable T-Sig analyses for actionable provenance analysis.
@@ 3.1 Video Transformer
AI-generated videos inherently exhibit domain gap [8, 6, 28], which is critical to address since the aim of SAGA is to be used in-the-wild. To enhance robustness, we extract potentially domain-agnostic features by leveraging a powerful visual encoder pretrained on web-scale image-text data. Given a video instance $x_{k}\in\mathcal{X}$, we process each frame $g_{m}$ (resized to a fixed resolution) using the frozen pretrained encoder. This produces a tokenized embedding $z_{m}\in\mathbb{R}^{l_{t}\times d_{t}}$ for each frame, where $m\in{1,2,\ldots,L}$ with $L$ denoting the number of frames per video. The dimension of these embeddings is determined by the chosen encoder, where $l_{t}$ is the number of tokens per frame and $d_{t}$ is the token feature dimension. The embeddings for all frames in $x_{k}$ are concatenated in temporal order, resulting in a video-level representation $\zeta_{k}\in\mathbb{R}^{L\times l_{t}\times d_{t}}$, which serves as input to our trainable video transformer. The resulting set of encoded videos is thus represented as $\mathcal{Z}={\zeta_{k}\mid x_{k}\in\mathcal{X}}$. The SAGA’s video-transformer model employs a multi-head self-attention (MHSA) transformer architecture [48] ($\theta$) tailored for video attribution to obtain $\phi_{k}=\theta(\zeta_{k})$. It processes sequences of frame embeddings, effectively capturing temporal dependencies for robust video-level predictions.
@@
  • A two-layer feed-forward network (MLP) with GELU activation [19] for non-linear transformations.
@@ 3.2 Contrastive Objective
Semi-hard negatives, which are most commonly used in the literature [39, 18, 55], are those that are further from the anchor than the positive, but within the margin as,
@@
Table 5: SOTA and SAGA comparison on DeMamba [7] for in-domain and DVF [42] for cross-domain evaluation. Best and second-best performances are highlighted.
@@
DeMamba [7] TALL [54] 87.91% 88.52% 88.42%
F3Net [34] 88.73% 81.88% 86.04%
NPR [45] 82.45% 84.08% 83.45%
STIL [16] 87.12% 82.22% 85.35%
MINTIME-CLIP-B [7] 91.55% 87.62% 89.98%
FTCN-CLIP-B [7] 92.21% 86.18% 89.67%
CLIP-B-PT [7] 44.83% 81.74% 41.82%
DeMamba-CLIP-PT [7] 79.97% 78.86% 79.98%
XCLIP-B-PT [7] 61.29% 81.93% 65.83%
DeMamba-XCLIP-PT [7] 76.38% 83.59% 79.31%
XCLIP-B-FT [7] 86.77% 84.41% 86.07%
@@
DVF [42] CNNDet [50] - - 78.20%
DIRE [51] - - 62.10%
Raising [13] - - 67.00%
UNI-FD [30] - - 74.10%
F3Net [34] - - 81.30%
ViViT [3] - - 79.10%
TALL [54] - - 69.50%
TS2-Net [27] - - 72.10%
DE-FAKE [40] - - 72.10%
HifiNet [17] - - 84.30%
DVF [42] - - 92.00%
@@ 4 Experiments
Datasets: The training is performed on the DeMamba [7] dataset, with 19 different AI video generators and 1M real videos. More details can be found in the supplementary material, including how we define different attribution levels. Additionally, we use the DVF dataset [42] covering 8 video generators for cross-data evaluations. Implementation details are provided in the supplementary.
Real/Fake Detection Results: We first comprehensively evaluate the performance of SAGA in binary classification in Table 4, conducting both in-domain and cross-generator evaluations. For cross-generator analysis, we train on the train split of DeMamba and evaluate on the val split, as well as train on generators from a specific team or SD version backbone or generation task and test on the remaining generators. The results demonstrate that, for the BIN-L task, SAGA achieves robust authenticity verification across diverse and previously unseen data sources, exhibiting strong generalization and minimal sensitivity to domain shifts. To rigorously benchmark SAGA against existing state-of-the-art (SOTA) binary detectors, we conducted evaluations on both the DeMamba [7] and DVF [42] datasets, as presented in Table 5. Notably, all competing SOTA methods on the DVF dataset [42] are trained and tested within the DVF dataset, following standard in-domain protocols. In contrast, SAGA is trained exclusively on DeMamba and evaluated directly on DVF, constituting a challenging cross-dataset generalization scenario. Despite this, SAGA significantly outperforms the SOTA baselines, highlighting its superior robustness and generalizability for the BIN-L task across diverse data distributions.
@@
t-SNE Analysis: To further interpret the representations learned by SAGA, we conduct a t-SNE [47] and T-Sig analyses of the feature embeddings produced by the model under different attribution settings. Specifically, we visualize the embedding outputs of the last ($6^{th}$) encoder of $\theta$ for videos from the validation set when the model is trained for all attribution levels. First, Fig. 4 (a) and (b) with TASK-L attribution results show that embeddings for real, T2V, and I2V samples form clearly separable clusters, demonstrating effective discrimination among these broad categories. However, when the same embeddings are colored by generator, substantial overlap is observed among most generators, with only a few, such as MorphStudio [43] and SVD [5], forming distinct clusters. This suggests that while SAGA is highly effective at coarse-grained attribution, it does not inherently separate individual generators at this level. Fig. 4 (c) shows the t-SNE plot for the SAGA model trained for BIN-L classification. Here, all generators except Pika [23] collapse into a single “fake” cluster, indicating that the model learns to aggregate all synthetic sources together for the binary task, with minimal separation among generators.
Fig. 4 (d) visualizes embeddings from a SAGA model trained on 10 specific generators from the DeMamba train set. The seen generators form distinct clusters, and notably, several unseen generators such as Hotshot [29], Show_1 [58], and MorphStudio [43] also appear as separable clusters. This indicates that the model, even trained on a subset of generators, can recognize distributional differences and cluster unseen sources, highlighting its potential for generalization in open-set scenarios. For the SD-L and TEAM-L models, t-SNE projections (Fig. 4 (e) - (h)) reveal that the learned representations not only cluster according to the supervised SD version or team labels, but also often separate individual generators within each group. This indicates that SAGA captures fine-grained differences between generators, even when supervision is provided only at a coarser level. This level of separation indicates that the model is sensitive to subtle distributional differences introduced by specific generator architectures or research teams, enabling it to infer whether an unknown generator shares an SD backbone or team affiliation, or represents a completely novel source.